One of the sums in Algebra states the below.
 
3x/4 = 9
x= 9*4/3
x=36/3
x=12
 
How is 3x/4 = 9 become 9*4/3?
 
Please clarify

Grade:6

4 Answers

Arun
4 years ago
Dear Aditi
 
3x/4 = 9
 
when side change is done then quantity in division converts into multiply
hence 4 is multiplied and 3 is in division
 
x = 9*4 /3 = 36/3 = 12
 
Vikas TU
4 years ago
Swati
4 years ago
Dear student, 
The given equation is 3x/4 = 9
i.e., 3x = 9 * 4 (because when a number is in division at one side and you want to take it on the other side , it will change into multiplication )
x = 36 / 3 (Same reason as above because division and multiplication are inter convertible)
implies,
x = 12
 
Hope this will help you out with your difficulty.
 
With regards,
Swati.
Yashraj
3 years ago
cross multiply 4 and then divide both sides by three 
hope this will help you ….......................................